OpenStudy (anonymous):

Find the solution for each of the following: a. f'(x)=4x, f(1)=6 b. f''(x)=2, f'(2)=5, f(2)=3

OpenStudy (amistre64):

dy/dx = 4x dy = 4x dx (S) dy = (S) 4x dx y = 2x^2 + C 6 = 2(1)^2 + C 6 = 2 + C 4 = C y = 2x^2 + 4

OpenStudy (amistre64):

d2y/dx2 = 2 d(dy) = (2 dx) dx I got no idea with this one.....

OpenStudy (anonymous):

f'' (x) = 2 f' (x) = 2x+c --> 2(2)+c=5 --> c=1 f(x) = x^2+c --> 4+c = 3 --> c=-1

OpenStudy (amistre64):

yay suzi!!!

OpenStudy (anonymous):

for b I got c=-3

OpenStudy (anonymous):

on the first deriv.

OpenStudy (anonymous):

I meant for the derivitive...not the first. I'm sorry.

OpenStudy (amistre64):

f'(x) = 2x+1 f(x) = x^2 + x + c right?

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

yay amistre! LOL

OpenStudy (amistre64):

i got me moments :)

OpenStudy (anonymous):

hahaha! It's okay! I have them too which is why I refer to this website! :)